A man walks 7km  in hours and 2km in 1 hour in the same direction. What is the man's averaged velocity of the whole journey

Grade:11

1 Answers

Harshit Singh
3 years ago
Dear Student
Your wuestion is missing some data, let us assume it like
A man walks 7km in 1 hours and 2km in 1 hour in the same direction. What is the man's averaged velocity of the whole journey

Total distance travelled = 7+2=9 km
Total time taken = 2 hours

Average Velocity = Total Distance/ Total time
= 9/2
= 4.5 kmph



Thanks